Main Street Ice Cream Company uses a plantwide allocation method to allocate overhead based on direct labor-hours at a rate of $3 per labor-hour. Strawberry and vanilla flavors are produced in Department SV. Chocolate is produced in Department C. Sven manages Department SV and Charlene manages Department C. The product costs (per thousand gallons) follow.

Strawberry Vanilla Chocolate Direct labor (per 1,000 gallons) $758 $833 $1,133
Raw materials (per 1,000 gallons) 808 508 608

Required: a. If the number of hours of labor per 1,000 gallons is 62 for strawberry, 66 for vanilla, and 100 for chocolate, compute the total cost of 1,000 gallons of each flavor using plantwide allocation.

b. Charlene's department uses older, outdated machines. She believes that her department is being allocated some of the overhead of Department SV, which recently bought state-of-the-art machines. After she requested that overhead costs be broken down by department, the following information was discovered: Department SV Department C Overhead $ 88,480 $ 43,778 Machine-hours 25,280 36,800 Labor-hours 25,280 18,550 Using machine-hours as the department allocation base for Department SV and labor-hours as the department allocation base for Department C, compute the allocation rate for each.

c. Compute the cost of 1,000 gallons of each flavor of ice cream using the department allocation rates computed in requirement (b) if the number of machine-hours for 1,000 gallons of each of the three flavors of ice cream are as follows: strawberry, 62; vanilla, 66; and chocolate, 158. Direct labor-hours by product remain the same as in requirement (a).
Answer #1

ANS 1(a) : Here we have to calculate the total cost Department wise and company wise as aggregate and in the question the total labor hours and the per hour cost of direct labor is give as $3 and the raw material is given so we will calculate as follows:

Ans1(b) : Here we need to calculate the allocation rate for each department which is

Overhead Cost of Department SV = $ 88480

Machine Hours =25280( It will be taken for the allocation rate of department SV)

Allocation rate of Department SV= $88,480 / 25280= $ 3.50.

Overhead Cost of Department C = $ 43,778

Labor Hours =18,550( It will be taken for the allocation rate of department C)

Allocation rate of Department C= $ 43,778 / 18,550= $ 2.36.

Ans 1( c) : Here we have to calculate the total cost Department wise and company wise as aggregate and in the question the total labor hours and the per hour cost of direct labor is give as $3 and as well as the machine hour rate of

SV machine cost= $ 3.5

C machine Cost = $ 2.36

Its calculation is as follows